# Weighted algebraic topology, II

### 05 May 2026 v1

@@ 0.6. Notation and terminology
Marginal remarks and elementary proofs are written in small characters.
@@ 0.7. Acknowledgements
I am indebted to my colleagues and friends Ettore Carletti, Fernando Gliozzi and Nicola Pinamonti, for long discussions on the present subject.
@@ 1. Values and weights
We begin by examining the extended real line $\overline{\mathbb{R}}=[{\overline{\infty}},\infty]$ and the extended half-line $\overline{\mathbb{R}}^{+}=[0,\infty]$, containing the values of the relative or positive metrics dealt with in the next section. The terms ‘relative’ and ‘positive’ will generally refer to these ordered sets.
As in Lawvere’s article [Lw], based on the extended half-line, lattice terminology will refer to the natural order of $\overline{\mathbb{R}}$ and $\overline{\mathbb{R}}^{+}$, while categorical terms will refer to the direction $\lambda\geqslant\mu$ of morphisms. This (unfortunate!) reversion is required by using numbers as values of metrics; for instance, the inequality $d\geqslant d^{\prime}$ between two metrics on the same set $X$ gives a ‘comparison map’ $(X,d)\to(X,d^{\prime})$ that lifts the identity.
1.1. Values, or relative weights
(a) The relative metrics we shall introduce take values in the extended line $\overline{\mathbb{R}}=[{\overline{\infty}},\infty]$, or – more precisely – in a structure $\bf w$} which can be analysed in two ways, based either on lattices and quantales [Mu, Ro, Ye] or on monoidal closed categories [EK, Ma, Ky]. A reader can focus on one of these aspects, but both are important: many points are more concrete in the terminology of lattices, but better motivated by category theory. (Some verifications are deferred to point (b).)
@@ 1.1. Values, or relative weights
(c) Joins and meets in $\bf w$} give ‘all’ limits and colimits in our category, that is, products and sums:
@@
The internal hom-object $\hom(\mu,\nu)\,$ is given by an extended difference, also written as $\nu-\mu$, and characterised as follows:
@@
The notation $\overline{\infty}$} allows us to distinguish the extended sum $\infty+{\overline{\infty}}=\infty$ and the extended difference $\infty-\infty={\overline{\infty}}$.
@@ 1.2. Positive values, or weights
(a) The complete lattice ${\bf w}^{+}=\overline{\mathbb{R}}^{+}=[0,\infty]$ of extended positive numbers, also called positive values, or weights, inherits a monoidal structure $\lambda+\mu$ which is also closed (in its own right). It was introduced by Lawvere [Lw], in 1974 (the original notation is $\bf R$}). Weight 0 should be viewed as ‘free’, and weight $\infty$ as ‘unreachable, unaffordable, or forbidden’. Metrics with values in $\bf w$$^{+}$ will be reviewed in 2.4.
The internal hom, written here as $\hom^{+}(\mu,\nu)$, is given by the truncated difference $(\nu-\mu){\,{\scriptstyle{\vee}}\,}0$:
@@
Truncated difference is written $\nu-\mu$ in [Lw], a notation introduced above in a different sense.
@@ 1.3. Multiplicative weights
The exponential and logarithm functions give an isomorphism (of monoidal closed categories):
@@
2. Relative and positive metrics
Lawvere’s generalised metric spaces [Lw] have been used in [G3, G4] as a first setting for weighted algebraic topology. Here we deal with the relative case.
@@ 2. Relative and positive metrics
2.1. Relative directed metrics
(a) A relative (directed) metric space $X$, or $\mu$-metric space, or $\mu$-space, will be a set $X$ equipped with a $\mu$-metric, that is, a mapping $\mu\colon X\!\times\!X\to{\bf w}$ satisfying the axioms
(21) $\mu(x,y)+\mu(y,z)\geqslant\mu(x,z),\qquad\;\;0\geqslant\mu(x,x)\qquad(x,y,z\in X).$
We shall generally write this structure $(X,\mu)$ as $X$, and the underlying set as $|X|$; the mapping $\mu$ is written as $\mu_{X}$ when useful.
The second condition above can be rewritten as: $\mu(x,x)=0$ or $\overline{\infty}$}, since the properties $0\geqslant\mu(x,x)$ and $\mu(x,x)+\mu(x,x)\geqslant\mu(x,x)$ are only satisfied by these extended values.
The value $\overline{\infty}$} is rather exceptional in $\mu$-spaces, as shown in 3.4(c), but excluding it would make bigger problems: a chaotic $\mu$-metric only takes values $\overline{\infty}$}, see 3.4(b).
(b) Formally, a $\mu$-space amounts to a small category $X$ enriched over the symmetric monoidal closed category $\bf w$} considered above, with a hom-value $X(x,y)$ in $\overline{\infty}$$\infty$], written as $\mu(x,y)$. This extended value will be viewed as the cost of the transition from $x$ to $y$, while $\gamma(x,y)=-\mu(x,y)$ represents the corresponding gain, or profit. (The function $\gamma$ is a ‘relative antimetric’, see 7.1.)
(c) The standard examples are related to the standard $\mu$-line $\mu\mathbb{R}$, with its cartesian and tensor powers
(22) $\mu\mathbb{R}=(\mathbb{R},\mu),\qquad\qquad\mu(x,y)=y-x.$
(d) $\mu\mathsf{Mtr}$ will denote the category of $\mu$-spaces, with cost-reducing maps $f\colon X\to Y$
(23) $\mu_{X}(x,x^{\prime})\geqslant\mu_{Y}(f(x),f(x^{\prime})),\qquad\text{for all }x,x^{\prime}\in X.$
Formally, $f$ is a w-enriched functor $X\to Y$, with hom-mapping $X(x,x^{\prime})$ $\to Y(f(x),f(x^{\prime}))$.
These mappings will also be called 1-Lipschitz maps of $\mu$-metric spaces, or $\mu$-maps. Isomorphisms in this category are isometric, and will be called isometric isomorphisms or 1-Lipschitz isomorphisms. More general ‘Lipschitz maps’ will be introduced in 2.5.
The reversor endofunctor $R\colon\mu\mathsf{Mtr}\to\mu\mathsf{Mtr}$ is based on the opposite $\mu$-metric space $R(X)=X^{{\rm op}}$, with the opposite $\mu$-metric, $\mu^{{\rm op}}(x,y)=\mu(y,x)$.
@@ 2.2. An overview of the main properties
The following properties of a $\mu$-metric space $X=(X,\mu)$ will be studied below.
(a) Positivity and finiteness. A positive $\mu$-metric, or $\delta$-metric, takes values in the coquantale $\bf w$$^{+}$; this case is reviewed in 2.4. A finite $\mu$-metric takes values in $\mathbb{R}$.
(b) Flatness. We say that a point $x\in X$ is regular (resp. flat) if $\mu(x,x)=0$ (resp. $\overline{\infty}$}). The space itself is regular (resp. flat) if all its points are. Every $\mu$-space $X$ has a flat part $X_{\overline{\infty}}$, formed by its flat points. A map of $\mu$-spaces can be restricted to the flat parts.
All important examples here are regular. Flat metrics include the chaotic ones, and are dealt with in 3.5.
(c) Linearity and invariance. There are two important classes of relative metrics. The standard $\mu$-line $\mu\mathbb{R}$, introduced in 2.1(c), belongs to both.
First, a linear $\mu$-metric will be a finite mapping $\mu\colon X\!\times\!X\to\mathbb{R}$ that satisfies the triangular identity: $\mu(x,y)+\mu(y,z)=\mu(x,z)$; in particular, $\mu(x,x)=0$ for all $x$. We shall see in 3.2 that a linear $\mu$-metric has the form $\Delta\varphi(x,y)=\varphi(y)-\varphi(x)$, for an arbitrary ‘potential’ function $\varphi\colon X\to\mathbb{R}$. Many examples related to diverse sciences, can be obtained in this way: for instance the gravitational potential determined by a distribution of maps or the electric potential determined by a distribution of charges, or the variation of any variable $\varphi$ (see 0.1 and 6.6).
Any chaotic metric satisfies the triangular identity; without the finiteness condition this is of little interest.
Secondly, a $\mu$-metric defined on an abelian group can be invariant under translations, or – equivalently – defined by a ‘relative norm’ (see 3.3). Most of the metrics we shall consider on the line and its cartesian or tensor products will be of this kind.
(d) Symmetry. A symmetric $\mu$-space has $\mu=\mu^{{\rm op}}$. They form a full subcategory $\mu\mathsf{Mtr}\subset\mu\mathsf{Mtr}$, which is reflective and coreflective, see 4.1 and 4.2. Proposition 3.6 will show that a symmetric $\mu$-space can be decomposed in a positive part and a flat part (which is empty, in the important examples).
More generally, according to a general definition in Directed Algebraic Topology ([G4], 1.2.1), the $\mu$-metric space $X$ is reversive if it is isometrically isomorphic to its opposite. This is the case of $\mu\mathbb{R}$ and its cartesian or tensor products.
@@ 2.3. Comparisons, limits and colimits
(a) The notation $X\geqslant X^{\prime}$ will mean that these $\mu$-spaces have the same underlying set and $\mu_{X}\geqslant\mu_{X^{\prime}}$, or equivalently that the identity of the underlying set is a $\mu$-map $X\to X^{\prime}$, called a comparison. We also say that $\mu_{X}$ is finer than $\mu_{X^{\prime}}$.
The $\mu$-metrics on a given set form a complete lattice: given a family of them $(\mu_{i})$, their pointwise join $\mu(x,y)={\,{\vee}\,}\mu_{i}(x,y)$ is easily seen to be a $\mu$-metric.
The category $\mu\mathsf{Mtr}$ has all limits (resp. colimits), computed as in $\mathsf{Set}$ and equipped with the least (resp. greatest) $\mu$-metric that makes the structural mappings into $\mu$-maps. We write the cartesian power $(\mu\mathbb{R})^{n}$ as $\mu\mathbb{R}^{n}$.
(b) A product $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}_{i}X_{i}\,$ has the $l_{\infty}$-type $\mu$-metric
(24) $\mu({\underline{x}},{\underline{y}})={\,{\vee}\,}\mu_{i}(x_{i},y_{i})\qquad\qquad({\underline{x}}=(x_{i}),\;{\underline{y}}=(y_{i})),$
and the terminal object is the $\mu$-singleton $\mu\top$, equipped with the least $\mu$-metric
(25) $\mu\top=\{*\},\qquad\mu(*,*)={\overline{\infty}}\qquad\qquad\text{(terminal object).}$
Equalisers have the restricted $\mu$-metric.
(c) A sum $\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,X_{i}$ has the obvious $\mu$-metric (and the initial object is $\varnothing$):
(26) $\mu((x,i),(y,i))=\mu_{i}(x,y),\qquad\mu((x,i),(y,j))=\infty\qquad(i\neq j).$
Coequalisers have the $\mu$-metric induced on the coequaliser in $\mathsf{Set}$, a quotient set.
Explicitly, the quotient $X/R$ of a $\mu$-metric space has the greatest $\mu$-metric making the canonical projection a $\mu$-map
(27) $\begin{array}[]{c}\mu(\xi,\eta)={\rm inf}_{\underline{x}}\;\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\mu(x_{2j-1},x_{2j}),\\[6.0pt] {\underline{x}}=(x_{1},...,x_{2n})\in X^{2n},\qquad\;\;x_{1}\in\xi,\;\;x_{2n}\in\eta,\\[4.0pt] x_{2j}R\,x_{2j+1}\;\text{ for }\;j=1,...,n-1.\end{array}$
@@ 2.4. Positive directed metrics
We recall the basic points of the generalised metrics introduced in [Lw]. They were used as a first setting for Weighted Algebraic Topology in [G3, G4].
(a) Following the terminology that we used in [G3, G4], a $\delta$-metric space $X$, or $\delta$-space, or positive $\mu$-space, is a set $X$ equipped with a $\delta$-metric $\delta\colon X\!\times\!X\to[0,\infty]$, also called a distance, satisfying the axioms
@@
(b) $\delta\mathsf{Mtr}$ denotes the full subcategory of $\mu\mathsf{Mtr}$ of $\delta$-metric spaces. Their cost-reducing maps are also called 1-Lipschitz maps of $\delta$-spaces, or $\delta$-maps.
@@
All non-empty products, all equalisers and all colimits are calculated as we have seen in $\mu\mathsf{Mtr}$.
(c) The embedding $\delta\mathsf{Mtr}\to\mu\mathsf{Mtr}$ has no left adjoint, as it does not preserve the terminal. The coreflector comes from the functor $(\;)^{+}\colon{\bf w}\to{\bf w}^{+}$ of (19), which truncates at 0 the negative values of a $\mu$-metric (yielding the smallest $\delta$-metric greater than the original $\mu$-metric)
(30) $\begin{array}[]{ccc}(\;\;)^{+}\colon\mu\mathsf{Mtr}\to\delta\mathsf{Mtr}&&\qquad\qquad\qquad\qquad\qquad\text{(coreflector),}\\[5.0pt] Y\mapsto Y^{+},&&\mu^{+}(y,y^{\prime})=(\mu(y,y^{\prime}))^{+}=\mu(y,y^{\prime}){\,{\scriptstyle{\vee}}\,}0,\\[5.0pt] Y^{+}\>\cong\>Y\!\times\!\delta\top,&&y\mapsto(y,*).\end{array}$
The absolute value does not give a functor from $\mu\mathsf{Mtr}$ to $\delta\mathsf{Mtr}$: for instance, it cannot work on the $\mu$-singleton, nor on any $\mu$-space having some flat point. But it does act on linear $\mu$-metrics, see 5.2(c).
@@
A symmetric $\delta$-metric $\delta=\delta^{{\rm op}}$ is the same as an écart in Bourbaki [Bk]. We shall denote by $\delta\mathsf{Mtr}$ the full subcategory of $\delta\mathsf{Mtr}$ of symmetric $\delta$-metric spaces. (This category was written $\mathsf{Mtr}$ in [G3, G4], but here we want to be consistent with the parallel case $\mu\mathsf{Mtr}$.)
(e) Structures on the line and the interval. The basic example, as in [G3, G4], is the standard $\delta$-line $\delta\mathbb{R}$, with $\delta$-metric:
@@
A different structure, used in [Lw], is also of interest: the $\delta_{0}$-line $\delta_{0}\mathbb{R}$
(32) $\delta_{0}\mathbb{R}=(\mu\mathbb{R})^{+}=(\mathbb{R},\delta_{0}),\qquad\delta_{0}(x,y)=(y-x){\,{\scriptstyle{\vee}}\,}0,$

### 13 May 2026 v2

@@ 0.7. Notation and terminology
Marginal remarks and elementary proofs are written in small characters. For the sake of brevity, some technical points are referred to the preprint version of this article [G5].
@@ 0.8. Acknowledgements
I am indebted to my colleagues Ettore Carletti and Fernando Gliozzi, for long discussions on the present subject.
@@ 1. Values and weights
We begin by examining the extended real line $\overline{\mathbb{R}}=[{\overline{\infty}},\infty]$ and the extended half-line $\overline{\mathbb{R}}^{+}=[0,\infty]$, containing the values of the real or positive metrics dealt with in the next section. The terms ‘real’ and ‘positive’ will generally refer to these ordered sets.
As in Lawvere’s article [L1], based on the extended half-line, lattice terminology will refer to the natural order of $\overline{\mathbb{R}}$ and $\overline{\mathbb{R}}^{+}$, while categorical terms will refer to the direction $\lambda\geqslant\mu$ of morphisms. This (unfortunate!) reversion is required by using numbers as values of metrics; for instance, the inequality $d\geqslant d^{\prime}$ between two metrics on the same set $X$ gives a ‘comparison map’ $(X,d)\to(X,d^{\prime})$ that lifts the identity. (Dealing with antimetrics would overcome this issue, but complicate the relationship with classical metrics.)
1.1. Values, or real weights
(a) The real metrics we shall introduce take values in the extended line $\overline{\mathbb{R}}=[{\overline{\infty}},\infty]$, or – more precisely – in a structure $\bf w$} which can be analysed in two ways, based either on lattices and quantales [Mu, Ro, Ye] or on monoidal closed categories [EK, Ma, Ky]. A reader can focus on one of these aspects, but both are important: many points are more concrete in the terminology of lattices, but better motivated by category theory. (Some verifications are deferred to point (b).)
@@ 1.1. Values, or real weights
(c) Joins and meets in $\bf w$} give ‘all’ limits and colimits in our category, which amount to products and sums
@@
The internal hom-object $\hom(\mu,\nu)\,$ is given by an extended difference, also written as $\nu-\mu$, and characterised as follows
@@
The notation $\overline{\infty}$} allows us to distinguish the extended sum $\infty+{\overline{\infty}}=\infty$ and the extended difference $\infty-\infty={\overline{\infty}}$. Note also that $\infty$ has no additive inverse in $\bf w$}.
@@ 1.2. Positive values, or weights
(a) The complete lattice ${\bf w}^{+}=\overline{\mathbb{R}}^{+}=[0,\infty]$ of extended positive numbers, also called positive values, or weights, inherits a monoidal structure $\lambda+\mu$ which is also closed (in its own right). It was introduced by Lawvere [L1], in 1974 (the original notation is $\bf R$}). Weight 0 should be viewed as ‘free’, and weight $\infty$ as ‘unreachable, unaffordable, or forbidden’. Metrics with values in $\bf w$$^{+}$ will be reviewed in 2.4.
The internal hom, written here as $\hom^{+}(\mu,\nu)$, is given by the truncated difference $(\nu-\mu){\,{\scriptstyle{\vee}}\,}0$
@@
Truncated difference is written $\nu-\mu$ in [L1], a notation introduced above in a different sense.
@@ 1.3. Multiplicative weights
The exponential and logarithm functions give an isomorphism (of monoidal closed categories)
@@
2. Real and positive metrics
Lawvere’s generalised metric spaces [L1] have been used in [G3, G4] as a first setting for weighted algebraic topology. Here we deal with the real case.
@@ 2. Real and positive metrics
2.1. Real valued metrics
(a) A real metric space $X$, or $\rho$-metric space, or $\rho$-space, will be a set $X$ equipped with a $\rho$-metric, that is, a mapping $\rho\colon X\!\times\!X\to{\bf w}$ satisfying the axioms
(21) $\rho(x,y)+\rho(y,z)\geqslant\rho(x,z),\qquad\;\;0\geqslant\rho(x,x)\qquad(x,y,z\in X).$
We shall generally write this structure $(X,\rho)$ as $X$, and the underlying set as $|X|$; the mapping $\rho$ is written as $\rho_{X}$ when useful.
The second condition above can be rewritten as: $\rho(x,x)=0$ or $\overline{\infty}$}, since the properties $0\geqslant\rho(x,x)$ and $\rho(x,x)+\rho(x,x)\geqslant\rho(x,x)$ are only satisfied by these extended values.
The value $\overline{\infty}$} is rather exceptional in $\rho$-spaces, as shown in 3.4(c), but excluding it would make bigger problems: a chaotic $\rho$-metric only takes the value $\overline{\infty}$}, see 3.4(b).
(b) Formally, a $\rho$-space amounts to a small category $X$ enriched over the symmetric monoidal closed category $\bf w$} considered above, with a hom-value $X(x,y)$ in $\overline{\infty}$$\infty$], written as $\rho(x,y)$. This extended value will be viewed as the cost of the transition from $x$ to $y$, while $\gamma(x,y)=-\rho(x,y)$ represents the corresponding gain, or profit. (The function $\gamma$ is a ‘real antimetric’, satisfying a reverse triangular inequality.)
(c) The standard examples are related to the standard $\rho$-line $\rho\mathbb{R}$, with its cartesian and tensor powers
(22) $\rho\mathbb{R}=(\mathbb{R},\rho),\qquad\qquad\rho(x,y)=y-x.$
(d) $\rho\mathsf{Mtr}$ will denote the category of $\rho$-spaces, with cost-reducing maps $f\colon X\to Y$
(23) $\rho_{X}(x,x^{\prime})\geqslant\rho_{Y}(f(x),f(x^{\prime})),\qquad\text{for all }x,x^{\prime}\in X.$
Formally, $f$ is a w-enriched functor $X\to Y$, with hom-map $X(x,x^{\prime})\to Y(f(x),f(x^{\prime}))$.
These mappings will also be called 1-Lipschitz maps of $\rho$-metric spaces, or $\rho$-maps. Isomorphisms in this category are isometric, and will be called isometric isomorphisms or 1-Lipschitz isomorphisms. More general ‘Lipschitz maps’ will be introduced in 2.5.
The reversor endofunctor $R\colon\rho\mathsf{Mtr}\to\rho\mathsf{Mtr}$ is based on the opposite $\rho$-metric space $R(X)=X^{{\rm op}}$, with the opposite $\rho$-metric, $\rho^{{\rm op}}(x,y)=\rho(y,x)$.
@@ 2.2. An overview of the main properties
The following properties of a $\rho$-metric space $X=(X,\rho)$ will be studied below.
(a) Positivity and finiteness. A positive $\rho$-metric, or $\delta$-metric, takes values in the coquantale $\bf w$$^{+}$; this case is reviewed in 2.4. A finite $\rho$-metric takes values in $\mathbb{R}$.
(b) Flatness. We say that a point $x\in X$ is regular (resp. flat) if $\rho(x,x)=0$ (resp. $\overline{\infty}$}). The space itself is regular (resp. flat) if all its points are. Every $\rho$-space $X$ has a flat part $X_{\overline{\infty}}$, formed by its flat points. A map of $\rho$-spaces can be restricted to the flat parts.
All important examples here are regular. Flat metrics include the chaotic ones. It is easy to see that a $\rho$-space $X$ is flat if and only if $\rho(x,x^{\prime})=\pm\infty$, for all $x,x^{\prime}\in X$.
Indeed, if $X$ is flat, $\rho(x,x^{\prime})\leqslant{\overline{\infty}}+\rho(x,x^{\prime})$, whence $\rho(x,x^{\prime})=\pm\infty$. The converse is also obvious, as the only possible values of $\rho(x,x)$ are 0 and $\overline{\infty}$}.
(c) Linearity and invariance. There are two important classes of real metrics. The standard $\rho$-line $\rho\mathbb{R}$, introduced in 2.1(c), belongs to both.
First, a linear $\rho$-metric will be a finite mapping $\rho\colon X\!\times\!X\to\mathbb{R}$ that satisfies the ‘additive property’ $\rho(x,y)+\rho(y,z)=\rho(x,z)$; in particular, $\rho(x,x)=0$ for all $x$. We shall see in Section 3.2 that a linear $\rho$-metric has the form $\Delta\Phi(x,y)=\Phi(y)-\Phi(x)$, for an arbitrary ‘potential’ function $\Phi\colon X\to\mathbb{R}$. Many examples related to diverse sciences, can be obtained in this way, using for instance the gravitational potential determined by a distribution of maps or the electric potential determined by a distribution of charges, or the variation of any variable $\Phi$ (see 0.2 and 6.6).
Any chaotic metric is additive, in this sense; without the finiteness condition this is of less interest.
Secondly, a $\rho$-metric defined on an abelian group can be invariant under translations, or – equivalently – defined by a ‘real norm’ (see 3.3). Most of the metrics we shall consider on the line and its cartesian or tensor products will be of this kind.
(d) Symmetry. A symmetric $\rho$-space has $\rho=\rho^{{\rm op}}$. They form a full subcategory $\rho\mathsf{Mtr}\subset\rho\mathsf{Mtr}$, which is reflective and coreflective, see 4.1 and 4.2.
Proposition 3.6 of [G5] shows that a symmetric $\rho$-space is the sum of a positive part and a flat part (which is empty, in the important examples).
More generally, according to a general definition in Directed Algebraic Topology ([G4], 1.2.1), the $\rho$-metric space $X$ is reversive if it is isometrically isomorphic to its opposite. This is the case of $\rho\mathbb{R}$ and its cartesian or tensor products.
@@ 2.3. Comparisons, limits and colimits
(a) The notation $X\geqslant X^{\prime}$ will mean that these $\rho$-spaces have the same underlying set and $\rho_{X}\geqslant\rho_{X^{\prime}}$, or equivalently that the identity of the underlying set is a $\rho$-map $X\to X^{\prime}$, called a comparison. We also say that $\rho_{X}$ is finer than $\rho_{X^{\prime}}$.
The $\rho$-metrics on a given set form a complete lattice: given a family of them $(\rho_{i})$, their pointwise join $\rho(x,y)={\,{\vee}\,}\rho_{i}(x,y)$ is easily seen to be a $\rho$-metric.
The category $\rho\mathsf{Mtr}$ has all limits (resp. colimits), computed as in $\mathsf{Set}$ and equipped with the least (resp. greatest) $\rho$-metric that makes the structural mappings into $\rho$-maps. We write the cartesian power $(\rho\mathbb{R})^{n}$ as $\rho\mathbb{R}^{n}$.
(b) A product $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}_{i}X_{i}\,$ has the $l_{\infty}$-type $\rho$-metric
(24) $\rho({\underline{x}},{\underline{y}})={\,{\vee}\,}\rho_{i}(x_{i},y_{i})\qquad\qquad({\underline{x}}=(x_{i}),\;{\underline{y}}=(y_{i})),$
and the terminal object is the $\rho$-singleton $\rho\top$, equipped with the least $\rho$-metric, the chaotic one
(25) $\rho\top=\{*\},\qquad\rho(*,*)={\overline{\infty}}\qquad\qquad\text{(terminal object).}$
Equalisers have the restricted $\rho$-metric.
(c) A sum $\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,X_{i}$ has the obvious $\rho$-metric (and the initial object is $\varnothing$)
(26) $\rho((x,i),(y,i))=\rho_{i}(x,y),\qquad\rho((x,i),(y,j))=\infty\qquad(i\neq j).$
Coequalisers have the $\rho$-metric induced on the coequaliser in $\mathsf{Set}$, a quotient set.
Explicitly, the quotient $X/R$ of a $\rho$-metric space has the greatest $\rho$-metric making the canonical projection a $\rho$-map
(27) $\begin{array}[]{c}\rho(\xi,\eta)={\rm inf}_{\underline{x}}\;\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho(x_{2j-1},x_{2j}),\\[6.0pt] {\underline{x}}=(x_{1},...,x_{2n})\in X^{2n},\qquad\;\;x_{1}\in\xi,\;\;x_{2n}\in\eta,\\[4.0pt] x_{2j}R\,x_{2j+1}\;\text{ for }\;j=1,...,n-1.\end{array}$
@@ 2.4. Positive directed metrics
We recall the basic points of the generalised metrics introduced in [L1]. They were used as a first setting for Weighted Algebraic Topology in [G3, G4].
(a) Following the terminology that we used in [G3, G4], a $\delta$-metric space $X$, or $\delta$-space, or positive $\rho$-space, is a set $X$ equipped with a $\delta$-metric $\delta\colon X\!\times\!X\to[0,\infty]$, also called a distance, satisfying the axioms
@@
(b) $\delta\mathsf{Mtr}$ denotes the full subcategory of $\rho\mathsf{Mtr}$ of $\delta$-metric spaces. Their (cost-reducing) maps are also called 1-Lipschitz maps of $\delta$-spaces, or $\delta$-maps.
@@
All non-empty products, all equalisers and all colimits are calculated as we have seen in $\rho\mathsf{Mtr}$.
(c) The embedding $\delta\mathsf{Mtr}\to\rho\mathsf{Mtr}$ has no left adjoint, as it does not preserve the terminal. The coreflector comes from the functor $(\;)^{+}\colon{\bf w}\to{\bf w}^{+}$ of (19), which truncates at 0 the negative values of a $\rho$-metric (yielding the smallest $\delta$-metric greater than the original $\rho$-metric)
(30) $\begin{array}[]{ccc}(\;\;)^{+}\colon\rho\mathsf{Mtr}\to\delta\mathsf{Mtr}&&\qquad\qquad\qquad\qquad\qquad\text{(coreflector),}\\[5.0pt] Y\mapsto Y^{+},&&\rho^{+}(y,y^{\prime})=(\rho(y,y^{\prime}))^{+}=\rho(y,y^{\prime}){\,{\scriptstyle{\vee}}\,}0,\\[5.0pt] Y^{+}\>\cong\>Y\!\times\!\delta\top,&&y\mapsto(y,*).\end{array}$
The absolute value does not give a functor from $\rho\mathsf{Mtr}$ to $\delta\mathsf{Mtr}$: for instance, it cannot work on the $\rho$-singleton, nor on any $\rho$-space having some flat point. But it does act on linear $\rho$-metrics, see 5.2(c).
@@
A symmetric $\delta$-metric $\delta=\delta^{{\rm op}}$ is the same as an écart in Bourbaki [Bk]. We shall denote by $\delta\mathsf{Mtr}$ the full subcategory of $\delta\mathsf{Mtr}$ of symmetric $\delta$-metric spaces. (This category was written $\mathsf{Mtr}$ in [G3, G4], but here we want to be consistent with the parallel case $\rho\mathsf{Mtr}$.)
(e) Structures on the line and the interval. The basic example, as in [G3, G4], is the standard $\delta$-line $\delta\mathbb{R}$, with $\delta$-metric
@@
A different structure, used in [L1], is also of interest, the $\delta_{0}$-line $\delta_{0}\mathbb{R}$
(32) $\delta_{0}\mathbb{R}=(\rho\mathbb{R})^{+}=(\mathbb{R},\delta_{0}),\qquad\delta_{0}(x,y)=(y-x){\,{\scriptstyle{\vee}}\,}0,$